OpenStudy (thomas5267):

Particle in a infinite potential well, which has a width of L. \[ E\psi (x)=\left[\frac{-\hbar ^2}{2m}\frac{d^2}{dx^2}+V(x)\right]\psi (x)\\ \psi _n (x)=\sqrt{\frac{2}{L}}\sin\left(\frac{\pi n x}{L}\right)\\ E_n=\frac{n^2\pi^2\hbar ^2}{2mL^2} \] The question is, what is the momentum of this particle?

OpenStudy (thomas5267):

I don't know whether it is valid to apply the momentum operator in this case. If I apply the momentum operator to the wavefunction, I get:\[ \begin{align*} &\hspace{1.5em}\mathbf{p}\psi_n(x)\\ &=-i\hbar\frac{d}{dx}\psi_n(x)\\ &=-ih\sqrt{\frac{2}{L}}\frac{\pi n}{L}\cos\left(\frac{\pi n x}{L}\right) \end{align*} \] Clearly you cannot have a imaginary momentum! Something must have gone wrong.

OpenStudy (michele_laino):

we can find the expection value of the momentum, or its value on the state represented by your wave function. Now that value is given by the subsequent formula: \[\Large p = \left\langle {{\psi ^*}} \right|\hat p\left| \psi \right\rangle = \int_0^L {{\psi ^*}} \left( { - i\hbar \frac{{d\psi }}{{dx}}} \right)dx\] I assumed that your problem is described by the subsequent drawing: 

OpenStudy (michele_laino):

I suggest to write your wave function as follows: \[\Large \begin{gathered} \psi \left( x \right) = A\left\{ {\exp \left( {i\Omega x} \right) - \exp \left( { - i\Omega x} \right)} \right\} = \hfill \\ \hfill \\ = 2iA\sin \left( {\Omega x} \right),\quad \left( {\Omega = \sqrt {\frac{{2mE}}{{{\hbar ^2}}}} } \right) \hfill \\ \end{gathered} \]

OpenStudy (thomas5267):

The drawing is correct. Is \(\Large p = \left\langle {{\psi ^*}} \right|\hat p\left| \psi \right\rangle\) the expectation value or the the momentum of the state? Furthermore, won't \(\hspace{1.5em} \Large\int_0^L {{\psi ^*}} \left( { - i\hbar \frac{{d\psi }}{{dx}}} \right)dx\\ \Large=\int_0^L\sqrt{\frac{2}{L}}\sin\left(\frac{\pi n x}{L}\right)\left(-ih\sqrt{\frac{2}{L}}\frac{\pi n}{L}\cos\left(\frac{\pi n x}{L}\right)\right)dx\) still be imaginary? One of the multiple choice question in my high school examination paper, which I finished a few days ago, asks for the the momentum of which a particle in a infinite potential well cannot have. However, I read somewhere that the momentum is not quantised for a particle in a infinite potential well. So I am trying to derive the momentum of a particle in a potential well and learn basic quantum mechanics on the way. Grateful if you could give me some advice.